You have $25,000 to invest in a stock portfolio. Your choices are Stock A with an expected return of 13 percent and Stock B with an expected return of 9 percent Required: (a) If your goal is to create a portfolio with an expected return of 11 percent, how much money will you invest in Stock B? (b)If your goal is to create a portfolio with an expected return of 13 percent, how much money will you invest in Stock A? can we label this portfolia as a well-diversified portfolio?
Answer #1

(a) Let money invested in stock B is x

Then, money invested in Stock A = $25,000 - x

Now, (25,000-x)0.13 + x*0.09 = 0.11*25,000

3,250 - 0.13x+0.09x = 2,750

X = $12,500

Hence, money invested in Stock B = $12,500

(B) Let money invested in Stock A be x

X*0.13 + (25,000-x)*0.09 = 0.13*25,000

X= $25,000

When, Money to be invested in Stock A = $25,000

This cannot be labelled as well diversified portfolio since all of the amount has been invested in the same stock. A well diversified portfolio is the one where the amount is invested in different securities.